A man walks at a pace of 3 miles per hour. After 2 hours he stops for a rest. Let t be time AFTER the rest when he continues on his walk. Let d be the TOTAL distance traveled since he started. Write an equation linking d and t.
Answer by ptaylor(1894)

Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
Distance travelled before the rest=3*2=6 mi
Distance travelled after the rest=3t
So, total distance (d)=dist before the rest+distance after the rest or
d=6+3t----------------------ans
Hope this helps---ptaylor